Your low-flow showerhead is delivering water at 1.1×10−4m3/s, about 1.8 gallons per minute. If this is the only water being used in your house, how fast is the water moving through your house's water supply line, which has a diameter of 0.021 m (about 3/4 of an inch)?
Answer #1

Solution) Q = 1.1×10^(-4) m^3/s = volume flow rate

Diameter d = 0.021 m

Radius r = d/2 = 0.021/2 = 0.0105 m

Q = Av

v = Q/A = Q/((pi)r^2)

v = (1.1×10^(-4))/((pi)(0.0105)^2)

v = 0.317 m/s